Calculate the fall in temperature when a gas initially at 72°C is expandedsuddenly to eight times its original volume. ( = 5/3)

Grade:11

1 Answers

Arun
6 years ago
Dear Vaibhav
 
T1 * V1(gamma-1) = T2* V2(gamma-1)
T2 = T1 (V1/V2)(gamma-1)
T2 = 345 *(x/8x)(5/3 -1)
on solving
T2 = 345 * (1/4) = 86.25 K
 
Regards
Arun (askIITians forum expert)